OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 7x + 3y = 30 -2x + 3y = 3

OpenStudy (anonymous):

...

OpenStudy (anonymous):

-2x+3y = 3 -2x = 3-3y x = (3-3y)/-2 x = (-3+3y)/2 7(-3+3y/2)+3y = 30 (-21+21y)/2+3y = 30 (-21+21y+6y)/2 = 30 -21+27y = 30*2 27y = 60+21 y = 81/27 y = 3 -2x+3*3 = 3 -2x+9 = 3 -2x = 3-9 x = -6/-2 x = 3 x=3,y=3